You are here: CLASSE Wiki>ACC/Bunch Web>BPMIssues (13 Dec 2005, MarkPalmer)Edit Attach

BPM Detector Issues

Ongoing issues with BPM 1W (Nov-Dec 2005)

BPM 1W Overview

  • This is a 6 button detector on the large round beampipe outside the IR
    • The top and bottom buttons are used for crossing measurements
    • Nominal 45 deg buttons are used for BPM
  • The top outside feedthrough was cross-threaded and physically distorted during the summer of 2003. This feedthrough had to be forced back into round (MAP/CRS) in order to be able to connect a cable to it. It is not clear whether any damage occurred to other feedthroughs. The distorted feedthrough continues to offer threading challenges in a tight space.
  • During recent operations 1W has regularly shown irregular readout - large position offsets and coupling errors.

December 6, 2005 Cable Check (MAP)

  • Cables were checked and connections tightened during Dec 6 access. No obvious problems found.

December 9, 2005 Orbit Response Characterization (MAP)

  • Characterized these errors on Dec 9 in preparation for down and possible replacment of BPM. NOTE degraded response of button 2!! Also note that the damaged feedthrough (top outer) corresponds to B4. Standard CESR button order is: 1=bot inner, 2=bot outer, 3=top inner, 4=top outer (this is not the same as the CBPM digital readout channel scheme).
    Buttons file DET B1 B2 B3 B4 TYPE Comments
    butns.109396 1 12402 8811 12643 13848 1 Baseline no seps
    butns.109397 1 5255 10092 5520 22686 1 HBL2 del = +1000
    butns.109398 1 10501 9640 10874 16021 1 HBL2 del = +200
    butns.109399 1 12300 8770 12532 13792 1 HBL2 0
    butns.109400 1 13093 9288 11864 12879 1 VBL2 del = +400
    butns.109401 1 11432 8207 13095 14601 1 VBL2 del = -400
    butns.109402 1 12215 8712 12481 13744 1 Back to 0
    butns.109403 1 15827 6706 15432 9727 1 HBL2 del = -400
    butns.109404 1 9170 9982 9586 17506 1 HBL2 del = +400
  • VBLIMING 2 lowered 400 CU:
    VBLIMING 2 lowered 400 CU
  • VBLIMING 2 raised 400 CU from nominal:
    VBLIMING 2 raised 400 CU from nominal
  • HBLIMING 2 lowered 400 CU from nominal:
    HBLIMING 2 lowered 400 CU from nominal
  • HBLIMING 2 raised 400 CU from nominal:
    HBLIMING 2 raised 400 CU from nominal

Dec 13, 2005 TDR Checkout and Cable Repairs (MAP/CRS)

  • Used the TDR to check response of all buttons. Disconnected cables at the 1W readout module on top of the west SCIR cryo platform.
    • Bottom Outside Button TDR showing poorly matched and distorted signal:
      Bottom Outside (bad signal) Button TDR
    • Bottom Inside Button TDR - good signal:
      Bottom Inside Button TDR
    • Top Inside Button TDR - good signal:
      Top Inside Button TDR:
    • Top Outside Button TDR - good signal:
      Top Outside Button TDR:
  • Since things clearly broken, decided to disconnect and swap cables to check whether problem could be in cable end. The problem was in the cable end - high impedance connection.
    • Bottom Outside Button after cable swap to use the nominal Top Outside cable:
      Bottom Outside Button after cable swap
  • Had Margee Carrier replace SMA end in situ. Due to cable length issues had to leave the Top Outside and Bottom Outside cables swapped. Once they checked out, we relabelled them so that all labels are now correct.
    • Bottom Outside cable after SMA replacement on Top Outside Button - cable relabelled and left connected to this button:
      Bottom Outside Cable after SMA replacement on Top Outside Button
    • Top Outside Cable on Bottom Outside Button - cable relabelled and left connected to this button:
      Top Outside Cable on Bottom Outside Button
  • 1805 - have a chance to re-time the CBPM module at 1W (changes in cable lengths).
    • Time in e+ and e- at 2mA and get new pedestal data
      • New configuration file: 252
    • Observe that button response from all 4 buttons is similar with SEPS OFF orbit
    • Did not want to hysterize machine with HBLIMING/VBLIIMING 2 orbit checks, so raise e- current to 4mA and verify that scaling of all 4 buttons is the same:
      • 2 mA raw data file: 2855
      • 4 mA raw data file: 2856
        Current B1 B2 B3 B4
        2 mA -113 -111 -109 -124
        4 mA -224 -220 -217 -249
      • So this problem appears fixed.
      • Note that we would still like to replace detector due to damaged B4 feedthrough!

Other Known Issues


  • This detector is on OLD CESR readout system. However, it is at the end of a long chain of only partially used detectors and quite often can go bad and not get noticed. Need to switch over to CBPM readout module.

BPM 6W/6W1

  • 6W is old system
  • 6W1 goes to CBPM test stand in room 101
  • Need to add 6W to CBPM system since can have similar issues as 3W.

Last Updated:  MarkPalmer - 13 Dec 2005

Topic revision: r1 - 13 Dec 2005, MarkPalmer
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding CLASSE Wiki? Send feedback